OpenStudy (anonymous):

evaluate 3^n+2-3^n/3^n

OpenStudy (anonymous):

is 8 the answer?

OpenStudy (anonymous):

3^n/3^n = 3^0 = 1

OpenStudy (anonymous):

3^n + 2 - 1

OpenStudy (anonymous):

3^n +2 - (-1)

OpenStudy (anonymous):

^n

OpenStudy (anonymous):

3^n +2 - (-1)^n is the answer

OpenStudy (anonymous):

how do you get last term?

OpenStudy (anonymous):

Since the −3 and 3 have the same exponent, they can be combined. Reduce the expression −3/3 by removing a factor of 3 from the numerator and denominator. Cancel the common factor of 3 in −3/3. Reduce the expression −3/3 by removing a factor of 3 from the numerator and denominator.